# Large-sample analysis of cost functionals for inference under the coalescent

### v1

@@ 5 Asymptotic analysis of importance sampling algorithms
For each $n\in\mathbb{N}$, we consider the truncated algorithms, starting at step 00 from a sample of the form $n\mathbf{y}_{0}^{(n)}$, satisfying Assumption 1.1, and stopping at step $k=\lfloor{tn}\rfloor{}$, for a fixed $t\in(0,1)$. To get an intuition about the extent of the truncation, consider the following. For large $n$, the starting sample size is $n\|\mathbf{y}^{(n)}_{0}\|_{1}\approx n\|\mathbf{y}_{0}\|_{1}=n$. After $\lfloor{tn}\rfloor{}$ steps, the sample size is reduced to $n\|\mathbf{Y}^{(n)}(\lfloor{tn}\rfloor{})\|_{1}$, where $\mathbf{Y}^{(n)}$ follows the proposal distribution. The latter is approximated by $n\|\mathbf{Y}(t)\|_{1}=n(1-t)$, as explained in the following proposition. This means that the truncated algorithms stop when the (large) sample size is reduced approximately by a factor $(1-t)$.
The sequence of Markov chains $\mathbf{Y}^{(n)}$, evolving backwards under the true distribution of Definition 2.2, converges to the deterministic trajectory $\mathbf{Y}$ (Theorem 3.3, and Favero and Hult (2024, Thm 2.1)). It is easy to see that the limit remains the same when $\mathbf{Y}^{(n)}$ evolves according to the GT or the SD proposal, after the importance sampling change of measure. This explains the approximation in Remark 5.1 and is more precisely stated in the following proposition for completeness.
@@
Let the scaled block-counting sequence of the coalescent $\mathbf{Y}^{(n)}$ evolve under the GT or SD proposal distribution. That is, $\mathbf{Y}^{(n)}$ is defined as in 2.4, but with backward transition probabilities given by the GT proposal (4.1) or by the SD proposal (4.2), rather than by Definition 2.2. Let $\mathbf{Z}^{(n)}=(C^{(n)},\mathbf{Y}^{(n)},\mathbf{M}^{(n)})$ be constructed from $\mathbf{Y}^{(n)}$ using Definitions 2.5 and 2.6. Then, under Assumptions 1.1 and 3.1, the convergence to the limiting process $Z$ of Theorem 3.3 is valid also for the sequence $\mathbf{Z}^{(n)}$ under the GT or the SD proposal distribution.
@@
$\frac{p(n\mathbf{Y}^{(n)}(\lfloor{tn}\rfloor{}))}{p(n\mathbf{y}_{0}^{(n)})}\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{1-d};\qquad C^{(n)}_{GT}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{d-1};\qquad C^{(n)}_{SD}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{d-1}.$
@@
$W^{(n)}_{GT}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}1;\qquad W^{(n)}_{SD}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}1;$
where ${\mathcal{D}}$ represents weak convergence, i.e. convergence in distribution.
@@
[…]
@@ 6.1 The finite alleles model
Figure 1 shows the empirical variance of importance weights in the GT and SD algorithms as a function of the remaining number of lineages. To generate it, independent replicate coalescent trees were initialised from the observed sample, and stopped as soon as they encounter a coalescence event. Once all replicates had been stopped, the variance of importance weights was recorded, simulation of all replicates was restarted, and the cycle of stopping replicates after each coalescence event was iterated until only one lineage remains in each replicate. To control runtimes, the GT scheme was run using the rejection control mechanism introduced in Section 5.2 of Griffiths and Tavaré (1994b), in which realisations with more than a given number of mutations are discarded. Throughout, we set the discard threshold to 1000.
@@
Figure 1: Log-variances of importance weights under the GT and SD proposals, measured by stopping replicates upon first hitting each fixed number of remaining lineages. Each figure is an averaged over 10 000 replicates.
The importance weight variances in both algorithms are plausibly converging towards 0 except in a region around the origin, where they spike very sharply. The convergence is especially rapid for the GT proposal. However, the relevant measure of algorithm performance is the maximal variance, which is 1-2 orders of magnitude higher for GT than SD across sample sizes, matching known results about the performance of these schemes (Stephens and Donnelly, 2000, Section 5).
@@
While it isn’t an informative indicator of overall algorithm performance, the low variance of weights for large samples evident in Figure 1 suggests that a small number of replicates could adequately represent the distribution of coalescent trees between the leaves and a low remaining sample size near the root. This would facilitate the allocation of more replicates to the sequential steps close to the root for a given computational budget. Allocating replicates into steps with high importance weight variance is known to be effective (Lee and Whiteley, 2018), but usually requires tuning via trial runs. Here this optimisation can be carried out a priori, at least heuristically.
@@
We tested this idea using the proposal $q_{SD}$ by initialising $\gamma=100$ independent replicate trees from the configuration of observed leaves $\mathbf{n}$, and simulating each until its number of remaining lineages first hit $\zeta<n$, whose value will be determined below. The resulting partially reconstructed trees were sampled with replacement until $\Gamma=10000$ were obtained, which were then independently propagated until the root.
@@
Our choice for the value of the threshold $\zeta$ is based on the ansatz that importance weights will begin to vary when the number of lineages has decreased due to coalescence by enough that mutations become commonplace. Before that point, proposed steps are predominantly coalescences between two lineages sharing a type, and the ordering of those events is unlikely to be important. The standard, untyped coalescent tree with $n$ leaves and mutation rate $\theta/2$ carries an average of $\theta\log(n)$ mutations when $n$ is large (Watterson, 1975). The probability that a given mutation occurs while there are between $\zeta$ and $n$ lineages in the tree is
@@
We simulated mutation rate likelihood estimators for a range of values using the SD proposal, $\chi=0.1$, and four different importance sampling schedules:
@@
The rationale for schedule 4 is that it simulates a constant number of replicates across all $n-1$ coalescence steps while expending approximately the same total computational effort as schedule 2. We neglect the random number of mutation steps when assessing computational effort because mutations are rare, and hence their contribution will be relatively small under the SD proposal. The approximate computational costs of executing all four schedules are depicted in Figure 2.
@@
Figure 2: Number of draws from the one-step proposal distribution $q_{SD}(\cdot|\cdot)$ for the four schedules with $\Gamma=10^{4}$, $\gamma=100$, $\theta=0.5$, and $\chi=0.1$. Note the log-scale on both axes.
@@
Figure 3: Performance of the four simulation schedules for various sample sizes based on independent simulations at points $\theta\in\{0.1,0.2,\ldots,0.9\}$. Standard errors were computed using the method of Chan and Lai (2013) for schedule 2, where realisations are not independent. The data-generating parameter is $\theta=0.5$.
@@
Figure 3 makes clear that the $10^{4}$ replicates of schedule 1 are needlessly expensive for accurate likelihood estimation when $n=50$. Schedule 3 with 100 replicates is by far the fastest but somewhat noisy. This is effect is exacerbated for $n=500$: schedule 3 remains the fastest but has large standard errors and does not appear smooth. Schedule 2 is also much faster than schedule 1, and nearly as accurate. Notably, it is both faster and slightly more accurate than schedule 4, so that the allocation of more replicates near the root at the cost of fewer replicates elsewhere is delivering a boost in accuracy. For $n=5000$ the same conclusion is even clearer: schedules 1 and 2 are virtually indistinguishable but the latter is faster by a factor of 24, while schedules 2 and 4 have noticeably larger standard errors.
So far we have focused on importance sampling without resampling. Figure 1 suggests that the variances of importance weights at intermediate times are not representative of their final variance, and begs the question of whether resampling based on importance weights is beneficial. It is well-known that, for the coalescent, resampling partially constructed replicates after a fixed number of simulation steps is harmful (Fearnhead, 2008). The standard remedy is so-called stopping-time resampling, in which partially reconstructed trees are stopped when the number of remaining lineages hits a given level, and resampling is performed once all replicates have been stopped (Chen et al., 2005; Jenkins, 2012). This schedule of resampling is an exact parallel of the method of stopping replicate simulation for representative importance weight variance calculation described above Figure 1. Figure 4 below makes clear that, for the standard coalescent and the SD proposal, resampling at these stopping times can also be harmful. For a less accurate proposal distribution, such as GT, stopping time resampling does dramatically improve inference (Chen et al., 2005, Section 6).
@@
Figure 4: A repeat of the top-left simulation in Figure 3 in which replicates were stopped whenever the number of lineages decreased. Once all replicates had stopped, systematic resampling (Chopin and Papaspiliopoulos, 2020, Section 9.6) was performed if the effective sample size (Kong et al., 1994) was less than 10% of the number of replicates.
@@ 6.2 The infinite sites model
The infinite sites model (ISM) is a more analytically and computationally tractable approximation of the site-by-site description of the finite alleles model. The genome of a lineage is associated with the unit interval $[0,1]$, which is also taken to be the type of the MRCA. Mutations occur along the branches of the coalescent tree with rate $\theta/2$, and each mutation is assigned to a uniformly sampled location along the genome. Mutations are inherited leaf-wards along the tree, so that the type of a sampled leaf is the list of mutations which occur on the branches connecting it to the MRCA. The list of mutations carried by an individual is referred to as its haplotype. The infinite sites approximation prohibits the same position mutating more than once, and is a good approximation when mutations are rare and the number of sites is large.
It is convenient to describe a sample of individuals from the infinite sites model as a triple $(\mathbf{S},\mathbf{n},\bm{\ell})$, where $\mathbf{S}$ is a matrix which lists observed haplotypes in its rows, with multiplicities given by $\mathbf{n}$, and where the location of each mutant site is listed in $\bm{\ell}$. If $h\leq n$ distinct haplotypes composed from a total of $r$ mutations are observed in a sample of $n$ individuals, then $\mathbf{S}$ is an $h\times r$ matrix with $S_{i,j}=1$ if haplotype $i$ carries mutation $j$, and 0 otherwise. The corresponding entry $n_{i}$ is the number of times haplotype $S_{i}=(S_{i,1},\ldots,S_{i,r})$ was observed, and $\ell_{j}\in[0,1]$ is the genomic location of the $j$th mutation.
@@
$p(\mathbf{S},\mathbf{n}^{\prime},\bm{\ell}^{\prime}|\mathbf{S},\mathbf{n},\bm{\ell})$
=$=\begin{cases}\frac{\|\mathbf{n}\|_{1}-1}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{i}}{\|\mathbf{n}\|_{1}}&\text{ if }(\mathbf{S}^{\prime},\mathbf{n}^{\prime},\bm{\ell}^{\prime})=(\mathbf{S},\mathbf{n}+\mathbf{e}_{i},\bm{\ell}),\quad i=1,\dots,h,\\ \frac{\theta}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{i}}{\|\mathbf{n}\|_{1}}&\text{ if }(\mathbf{S}^{\prime},\mathbf{n}^{\prime},\bm{\ell}^{\prime})=(E_{ij}\mathbf{S},a_{j}(\mathbf{n},1),a_{j}(\bm{\ell},x)),\quad i=1,\dots h,\;j=0,\ldots,r,\\ 0&\text{ otherwise},\end{cases}$
where $a_{j}(\mathbf{v},x)$ is the vector obtained from $\mathbf{v}$ by inserting the scalar $x$ between then $j$th and $(j+1)$th positions, and $E_{ij}$ is an operator which inserts a duplicate of row $i$ as the new last row of $\mathbf{S}$, and then inserts $\mathbf{e}_{h+1}$ as a new column in the $j$th position. The backward transition probabilities are intractable, similarly to the finite alleles case, and don’t depend on the labels $\bm{\ell}$ so we suppress them from the notation going forward, for the sake of readability.
@@
$q_{GT}(\mathbf{S}^{\prime},\mathbf{n}^{\prime}|\mathbf{S},\mathbf{n})$$\propto\begin{cases}(n_{j}-1)&\text{if }(\mathbf{S}^{\prime},\mathbf{n}^{\prime})=(\mathbf{S},\mathbf{n}-\mathbf{e}_{j})\text{ and }n_{j}\geq 2,\\ \theta(n_{j^{\prime}}+1)/\|\mathbf{n}\|_{1}&\text{if }n_{j}=1,j\in\mathcal{M},\text{ and }\exists\omega\;\&\;j^{\prime}\neq j:(S^{\prime})_{j^{\prime}}=S_{j}^{\omega},\\ \theta/\|\mathbf{n}\|_{1}&\text{if }n_{j}=1,j\in\mathcal{M},\text{ and }\exists\omega:(S^{\prime})_{j}=S_{j}^{\omega},\\ 0&\text{otherwise},\end{cases}$
@@
improving the scaling in both sample size and sequence length by a linear factor. However, the SD proposal is substantially faster at a cost of $O(h)$ per step, or
@@
Theorem 3.3 does not apply to the ISM. However, since the ISM is regarded as a good approximation to the finite alleles model for long sequences and rare mutations, it is instructive to examine whether similar conclusions about importance sampling proposal distributions hold. To that end, we applied all three ISM proposal distributions to the data set of Ward et al. (1991)—a common benchmark with $n=55$ samples and $r=18$ mutations. To assess scaling, we also simulated two synthetic data sets with respective sizes $n=550$ and $n=5500$ using $\theta=5.0$, which is the approximate maximum likelihood estimator from the Ward et al. (1991) data set. For HUW, we set the driving value of $\theta$ used to pre-compute the proposals for each data set equal to the Watterson estimator (Watterson, 1975), which takes respective values 3.93, 4.94, and 4.90 for the three data sets. The largest matrix took around 2 hours of computing time in serial, but the computation is trivial to parallelise and can be reused for any data set with size no greater than 5500 and for which 4.9 is an acceptable driving value for the mutation rate.
Figure 5: Log-variances of importance weights for the GT, SD, and HUW proposals, measured by stopping replicates upon first hitting each fixed number of remaining lineages. Each figure was obtained by averaging $10^{5}$ replicates.
@@
We tested these suggestions by simulating likelihood estimators independently for a range of values of $\theta$, using the four replicate schedules from Section 6.1. Figure 6 bears out both suggestions for the data set with $n=55$ samples: the results with resampling are considerably less noisy than those without, except for schedule 3 with only 1000 particles which has very high standard error. There is also very little difference between schedules 1, 2, and 4. Figure 7 shows that the same conclusions hold for a larger data set with $n=550$ samples. It also illustrates the difference in computational cost between the HUW and SD proposals, which was already evident in the per-replicate analyses in (6.2) and (6.3). The gains in accuracy with the HUW proposal do not seem to compensate for its higher cost.
Figure 6: Likelihood estimates for the $n=55$ data set from the HUW and SD proposals, simulated using the four schedules of replicates with $\gamma=10^{3}$ and $\Gamma=10^{5}$, independently for $\theta\in\{2,3,\ldots,10\}$. Replicates in the right column were resampled in the way described in the caption of Figure 4. Standard errors for schedule 2, and for every schedule with resampling, were computed using the unbiased method of Chan and Lai (2013).
Figure 7: Likelihood estimates for the $n=55$ data set from the HUW and SD proposals, simulated using the four schedules of replicates with $\gamma=2\times 10^{3}$ and $\Gamma=2\times 10^{5}$, independently for $\theta\in\{2,3,\ldots,10\}$. Replicates in the right column were resampled in the way described in the caption of Figure 4. Standard errors for schedule 2, and for every schedule with resampling, were computed using the unbiased method of Chan and Lai (2013).

### v2

@@ 5 Asymptotic analysis of importance sampling algorithms
For each $n\in\mathbb{N}$, we analyse truncated algorithms starting from a sample of the form $n\mathbf{y}_{0}^{(n)}$ satisfying Assumption 1.1, and stopping after $k=\lfloor{tn}\rfloor{}$ steps, for a fixed $t\in(0,1)$. To get an intuition about the extent of the truncation, consider the following. For large $n$, the starting sample size is $n\|\mathbf{y}^{(n)}_{0}\|_{1}\approx n\|\mathbf{y}_{0}\|_{1}=n$. After $\lfloor{tn}\rfloor{}$ steps, the sample size is reduced to approximately $n\|\mathbf{Y}^{(n)}(\lfloor{tn}\rfloor{})\|_{1}$, where $\mathbf{Y}^{(n)}$ follows the proposal distribution. The latter is approximately $n\|\mathbf{Y}(t)\|_{1}=n(1-t)$, as justified by Proposition 5.2 below. This means that the number of lineages remaining when the truncated algorithm is stopped after $\lfloor tn\rfloor$ steps is approximately n(1t)n(1-t). Essentially all practical algorithms are untruncated, corresponding to $t=1$, but the technical details of our asymptotics only hold for $t<1$. In Section 6, we show empirically that our results are robust to this discrepancy, in that none of the simulations we carry out use truncated algorithms.
The sequence of Markov chains $\mathbf{Y}^{(n)}$, evolving backwards under the true distribution of Definition 2.2, converges to the deterministic trajectory $\mathbf{Y}$ (see Theorem 3.3, and Favero and Hult (2024, Thm 2.1)). It is easy to see that the limit remains the same when $\mathbf{Y}^{(n)}$ evolves according to the GT or the SD proposal, after the importance sampling change of measure. This explains the approximation in Remark 5.1 and is more precisely stated in the following proposition for completeness.
@@
Let the scaled block-counting sequence of the coalescent $\mathbf{Y}^{(n)}$ evolve under the GT or SD proposal distribution. That is, $\mathbf{Y}^{(n)}$ is defined as in 2.4, but with backward transition probabilities given by the GT proposal (4.1) or by the SD proposal (4.2), rather than by Definition 2.2. Let $\mathbf{Z}^{(n)}=(C^{(n)},\mathbf{Y}^{(n)},\mathbf{M}^{(n)})$ be constructed from $\mathbf{Y}^{(n)}$ using Definitions 2.5 and 2.6. Then, under Assumptions 1.1 and 3.1, the convergence to the limiting process $\mathbf{Z}$ of Theorem 3.3 is valid also for the sequence $\mathbf{Z}^{(n)}$ under the GT or the SD proposal distribution.
@@
$\frac{p(n\mathbf{Y}^{(n)}(\lfloor{tn}\rfloor{}))}{p(n\mathbf{y}_{0}^{(n)})}\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{1-d};\hskip 18.49988ptC^{(n)}_{GT}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{d-1};\hskip 18.49988ptC^{(n)}_{SD}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{d-1}.$
@@
$W^{(n)}_{GT}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}1;\hskip 18.49988ptW^{(n)}_{SD}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}1;$
where $\xrightarrow[]{\mathcal{D}}$ represents weak convergence, i.e. convergence in distribution.
@@
$W^{(n)}_{*}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}(1-t)^{1-d}\exp\left\{\int_{0}^{t}\sum_{i=1}^{d}y_{0,i}\int_{0}^{t}a^{*}_{i}\left(\mathbf{y}_{0}(1-u)\right)du\right\}=1.$
@@
Remark 5.4 gives a criterion on importance sampling proposal distributions under which importance weights do not degenerate with large sample size. However, by itself the lack of degeneracy indicates very little about the quality of the resulting algorithm in practice. This is evident because both the GT and SD proposals satisfying 5.4, yet the latter yields a far superior algorithm. In our opinion, the practical utility of Theorem 5.3 and Remark 5.4 is in informing the design of particle allocation schedules and resampling schemes for a given proposal distribution, as we show by simulation in Section 6 below.
@@ 6.1 The finite alleles model
Figure 1 shows empirical normalised second moments of importance weights $W_{1},\ldots,W_{N}$ of $N$ simulated replicates in the GT and SD algorithms as a function of the remaining number of lineages. The empirical normalised second moment is related to the usual effective sample size (ESS) of Kong et al. (1994) via
$\frac{\frac{1}{N}\sum_{i=1}^{N}W_{i}^{2}}{\big(\frac{1}{N}\sum_{i=1}^{N}W_{i}\big)^{2}}=\frac{N}{\text{ESS}}.$ (6.1)
@@
To generate Figure 1, independent replicate coalescent trees were initialised from the observed sample, and stopped as soon as they encounter a coalescence event. Once all replicates had been stopped, the ESS was recorded, simulation of all replicates was restarted, and the cycle of stopping replicates after each coalescence event was iterated until only one lineage remains in each replicate. To control runtimes, the GT scheme was run using the rejection control mechanism introduced in Section 5.2 of Griffiths and Tavaré (1994b), in which realisations with more than a given number of mutations are discarded. Throughout, we set the discard threshold to 1000.
@@
Figure 1: Logarithms of normalised second moments of importance weights under the GT and SD proposals, measured by stopping replicates upon first hitting each fixed number of remaining lineages. Each figure is an average over 10 000 replicates.
@@
The importance weight variances in both algorithms are plausibly converging towards 0 as $n\to\infty$ except in a neighbourhood of the origin, where they spike sharply. The convergence is especially rapid for the GT proposal. Given that the SD proposal is known to be considerably more accurate than GT (Stephens and Donnelly, 2000, Section 5), the precise connection between these variance profiles and algorithm performance is not transparent. However, while it is not an informative indicator of overall algorithm performance, the low variance of weights for large samples evident in Figure 1 suggests that a small number of replicates could adequately represent the distribution of coalescent trees between the leaves and a low remaining number of lineages near the root. This would facilitate the allocation of more replicates to the sequential steps close to the root for a given computational budget. Allocating replicates into steps with high importance weight variance is known to be effective (Lee and Whiteley, 2018), but usually requires tuning via trial runs. Here this optimisation can be carried out a priori, at least heuristically.
@@
We tested this idea using the proposal $q_{SD}$ by initialising $\gamma=100$ independent replicate trees from the configuration of observed leaves $\mathbf{n}$, and simulating each until its number of remaining lineages first hit $\zeta<n$, whose value will be determined below. The resulting partially reconstructed trees were sampled with replacement until $\Gamma=10000$ replicates were obtained, which were then independently propagated until the root.
@@
Our choice for the value of the threshold $\zeta$ is based on the ansatz that importance weights will begin to vary when the number of lineages has decreased due to coalescence by enough that mutations become commonplace. Before that point, most proposed steps are coalescences between two lineages which share a type, and the ordering of those events is unlikely to be important. The standard, untyped coalescent tree with $n$ leaves and mutation rate $\theta/2$ carries an average of $\theta\log(n)$ mutations when $n$ is large (Watterson, 1975). The probability that a given mutation occurs while there are between $\zeta$ and $n$ lineages in the tree is
@@
We simulated sampling probability estimators for a range of mutation rates using the SD proposal, $\chi=0.1$, and four different importance sampling schedules:
@@
The rationale for schedule 4 is that it simulates a constant number of replicates across all n1n-1 coalescence steps while expending approximately the same total computational effort, measured as the total number of simulated coalescence steps, as schedule 2. We neglect the random number of mutation steps when assessing computational effort because their number is not predictable under the finite alleles model. However, mutations are rare under the SD proposal, and hence their contribution will be small with very high probability.
@@
Figure 2: Number of simulated coalescence steps from the one-step proposal distribution $q_{SD}(\cdot|\cdot)$ for the four schedules with $\Gamma=10^{4}$, $\gamma=100$, $\theta=0.5$, and $\chi=0.1$. Note the log-scale on both axes.
The approximate computational effort of executing all four schedules are depicted in Figure 2. The purpose of these schedules is to assess how the number of simulated replicates interacts with the sample size $n$ and the presence or absence of resampling, and whether or not the non-uniform allocation of computational effort in schedule 2 is beneficial. Both practical questions are motivated by Theorem 3.3, which suggests that importance sampling methods might stabilise for large sample sizes even for a fixed number of particles, and that consequently, it is advantageous to expend more computational effort near the root of the tree where the number of remaining lineages is low. This is in stark contrast to general importance sampling theory, which dictates that the number of replicates needs to grow exponentially in the dimension of the latent object (Chatterjee and Diaconis, 2018, Theorem 1.1). Here, the dimension of the latent ancestral tree is approximately linear in sample size.
@@
Figure 3: Performance of the four schedules with $\gamma=10^{2}$ and $\Gamma=10^{4}$ for various sample sizes based on independent simulations at points $\theta\in\{0.1,0.2,\ldots,0.9\}$. Standard errors were computed using the method of Chan and Lai (2013) for schedule 2, where replicates are not independent. The data-generating parameter is $\theta=0.5$. Each y-axis is multiplied by appropriate, large constant to aid visualisation, and small horizontal offsets have been artificially added to all four curves in each panel for visual clarity.
@@
Figure 3 shows simulated sampling distribution estimators for three different sample sizes using all four schedules and a range of values of the mutation rate $\theta$. Good performance of a schedule is indicated by a smooth graph with small standard errors. In line with the prediction of Theorem 3.3, there is no noticeable reduction in performance for larger sample sizes: $n=50$ and $n=5000$ show comparable performance with the same number of simulated replicates. It is also clear that the $10^{4}$ replicates of schedule 1 are needlessly expensive for accurate sampling probability estimation. Schedule 3 with 100 replicates is by far the fastest but somewhat noisy. Schedule 2 is also much faster than schedule 1, and nearly as accurate. Notably, it is both faster and slightly more accurate than schedule 4, so that the allocation of more replicates near the root at the cost of fewer replicates elsewhere is delivering a boost in accuracy. For $n=5000$, the most challenging case, schedules 1 and 2 are virtually indistinguishable but the latter is faster by a factor of 24, while schedules 3 and 4 have noticeably larger standard errors.
@@
Figure 4: A repeat of the simulation in Figure 3(a) in which replicates were stopped whenever the number of lineages decreased. Once all replicates had stopped, systematic resampling (Chopin and Papaspiliopoulos, 2020, Section 9.6) was performed if the effective sample size, ESS in (6.1), was less than 10% of the number of replicates. The y-axis is expressed in units of $10^{-13}$ to aid visualisation, and small horizontal offsets have been artificially added to all four curves for visual clarity.
So far we have focused on importance sampling without resampling. Figure 1 suggests that importance weights at intermediate times are poor predictors of their final values, and invites the question of whether resampling based on importance weights is beneficial. It is well-known that, for the coalescent, resampling partially constructed replicates after a fixed number of simulation steps is harmful (Fearnhead, 2008). The standard remedy is so-called stopping-time resampling, in which partially reconstructed trees are stopped when the number of remaining lineages hits a given level, and resampling is performed once all replicates have been stopped (Chen et al., 2005; Jenkins, 2012). This schedule of resampling is an exact parallel of the method of stopping replicate simulations for representative importance weight variance calculation described above Figure 1. Figure 4 makes clear that, for the standard coalescent and the SD proposal, resampling at these stopping times can also be harmful: sampling probability estimates are noisier and have larger standard errors than those in Figure 3(a), which depicts the same simulation without resampling. Schedules 3 and 4 in particular depict substantial Monte Carlo error, which results in incorrect estimates of both the mean and standard error. It is noteworthy that the method depicted in Figure 4 used a restrictive resampling threshold, $\text{ESS}<N/10$, which typically only triggered 2–5 times per run for $N=10^{4}$ particles in this example. Thus, even a small amount of resampling had a marked, harmful effect on estimator performance. Typical guidance on resampling for generic particle filtering and SMC is to resample when $\text{ESS}<N/2$ (Chopin and Papaspiliopoulos, 2020, Section 10.2). In this case the change in threshold would make little difference since, as shown in Figure 1, importance weight variance is low for most of the run before spiking rapidly at the end. For a less accurate proposal distribution such as GT, stopping time resampling does dramatically improve inference, and it has also been observed to be beneficial for the SD proposal under a microsatellite model of mutation (Chen et al., 2005, Section 6).
@@ 6.2 The infinite sites model
The infinite sites model (ISM) is a more analytically and computationally tractable approximation of the site-by-site description of the finite alleles model. The genome of a lineage is associated with the unit interval $[0,1]$, which is also taken to be the type of the MRCA. Mutations occur along the branches of the coalescent tree with rate $\theta/2$, and each mutation is assigned to a uniformly sampled location along the genome. Mutations are inherited along the tree towards the leaves, so that the type of a sampled leaf is the list of mutations which occur on the branches connecting it to the MRCA. The list of mutations carried by an individual is referred to as its haplotype. The infinite sites approximation prohibits the same position mutating more than once, and is a good approximation when mutations are rare and the number of sites is large.
[…]
@@
$p(\mathbf{S}^{\prime},\mathbf{n}^{\prime},\bm{\ell}^{\prime}|\mathbf{S},\mathbf{n},\bm{\ell})$
$=\begin{cases}\frac{\|\mathbf{n}\|_{1}-1}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{i}}{\|\mathbf{n}\|_{1}}&\text{ if }(\mathbf{S}^{\prime},\mathbf{n}^{\prime},\bm{\ell}^{\prime})=(\mathbf{S},\mathbf{n}+\mathbf{e}_{i},\bm{\ell}),\qquad i\in\{1,\dots,h\},\\ \frac{\theta}{\|\mathbf{n}\|_{1}-1+\theta}\frac{n_{i}}{\|\mathbf{n}\|_{1}}&\text{ if }(\mathbf{S}^{\prime},\mathbf{n}^{\prime},\bm{\ell}^{\prime})=(E_{ij}\mathbf{S},a_{j}(\mathbf{n},1),a_{j}(\bm{\ell},x)),\qquad i\in\{1,\dots h\},\;j\in\{0,\ldots,r\},\\ 0&\text{ otherwise},\end{cases}$
[…]
@@
[…]
@@
improving the scaling in both sample size and sequence length by a linear factor. However, the infinite sites SD proposal is substantially faster at a cost of $O(h)$ per step, or
@@
Theorem 3.3 does not apply to the ISM. However, since the ISM is regarded as a good approximation to the finite alleles model for long sequences and rare mutations, it is instructive to examine whether similar conclusions about importance sampling proposal distributions hold. To that end, we applied all three ISM proposal distributions to the data set of Ward et al. (1991)—a common benchmark with $n=55$ samples and $r=18$ mutations. To assess scaling, we also simulated two synthetic data sets with respective sizes $n=550$ and $n=5500$ using $\theta=5.0$, which is the approximate maximum likelihood estimator from the Ward et al. (1991) data set. For HUW, we set the driving value of $\theta$ used to pre-compute the proposals for each data set equal to the Watterson estimator (Watterson, 1975), which takes respective values 3.93, 4.94, and 4.90 for the three data sets. The largest matrix took around 2 hours of computing time in serial, but the computation is trivial to parallelise and can be reused for any data set with size no greater than 5500 and for which 4.9 is an acceptable driving value for the population-rescaled mutation rate.
Figure 5: Logarithms of normalised second moments of importance weights for the GT, SD, and HUW proposals, measured by stopping replicates upon first hitting each fixed number of remaining lineages. Each figure was obtained by averaging $10^{5}$ replicates. The results from Figure 1(a) and 1(b) are reproduced in dashed lines in panels 5(a) and 5(b) for ease of comparison.
@@
We tested these suggestions by simulating sampling distribution estimators independently for a range of values of $\theta$, using the four replicate schedules from Section 6.1. Figure 6 bears out both suggestions for the data set with $n=55$ samples: the results with resampling are considerably less noisy than those without, except for schedule 3 with only 1000 particles which has very high standard error. There is also very little difference between schedules 1, 2, and 4. Schedule 3 becomes erratic when resampling is applied, showing substantial Monte Carlo error in estimates of both the mean and standard error. Based on non-overlapping confidence intervals, schedules 1, 2, and 4 also appear to underestimate standard errors slightly when resampling is used.
Figure 6: Sampling probability estimates for the $n=55$ data set from the HUW and SD proposals, simulated using the four schedules of replicates with $\gamma=10^{3}$ and $\Gamma=10^{5}$, independently for $\theta\in\{2,3,\ldots,10\}$. Replicates in panels 6(b) and 6(d) were resampled in the way described in the caption of Figure 4. Standard errors for schedule 2, and for every schedule with resampling, were computed using the unbiased method of Chan and Lai (2013). Each y-axis is expressed in units of $10^{-20}$ to aid visualisation, and small horizontal offsets have been artificially added to all four curves in each panel for visual clarity.
Figure 7 shows that similar conclusions hold for a larger data set with $n=550$ samples. It also illustrates the difference in computational cost between the HUW and SD proposals, which was already evident in the per-replicate analyses in (6.3) and (6.4). The gains in accuracy with the HUW proposal do not seem to compensate for its higher cost. Comparing Figures 6 and 7 clearly shows that the performance of both the SD and HUW proposals deteriorates with increasing sample size, unlike in Figure 3 for the finite alleles model. None of the schedules appear fully reliable, with plenty of evidence of Monte Carlo error and ill-calibrated standard error estimates. Resampling appears to be beneficial, and indeed the means of schedules 1 and 2 nearly coincide when resampling takes place, though their standard error estimates still yield non-overlapping confidence intervals in some cases.
Figure 7: Sampling probability estimates for the $n=550$ data set from the HUW and SD proposals, simulated using the four schedules of replicates with $\gamma=2\times 10^{3}$ and $\Gamma=2\times 10^{5}$, independently for $\theta\in\{2,3,\ldots,10\}$. Replicates in panels 7(b) and 7(d) were resampled in the way described in the caption of Figure 4. Standard errors for schedule 2, and for every schedule with resampling, were computed using the unbiased method of Chan and Lai (2013). Eaxh y-axis is expressed in units of $10^{-46}$ to aid visualisation, and small horizontal offsets have been artificially added to all four curves in each panel for visual clarity.